OpenStudy (anonymous):

Fill in the missing number to complete the factorization: x^3+x^2-21x-45=(x+3)(x-5)(x+_________)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Is the constant term 3 @ganeshie8

ganeshie8 (ganeshie8):

constant term is not 3

OpenStudy (anonymous):

45=3*5=15 45/15=3

OpenStudy (anonymous):

AWW

ganeshie8 (ganeshie8):

your answer is 3, though

OpenStudy (anonymous):

I:

ganeshie8 (ganeshie8):

constant term on left hand side is -45

OpenStudy (anonymous):

-3 then

ganeshie8 (ganeshie8):

when you use the terms correctly, others will know that you know the concepts correctly

ganeshie8 (ganeshie8):

you need to solve : \[-45 = 3(-5)*c\]

ganeshie8 (ganeshie8):

\[45 = 15c \implies c = 3\]